OpenStudy (beggi9):

Particle moves along a straight line with the acceleration of a = 2 / sqrt (v) where v = dx/dt is the velocity. When t=2s then x = 64/3m with speed v = 16m/s. Determine the location and acceleration when t= 3s

OpenStudy (irishboy123):

break it down you have \(a = \dfrac{2}{\sqrt{v}}\) which means that \(\dfrac{dv}{dt} = \dfrac{2}{\sqrt{v}}\) So \(\int \sqrt{v} \, dv =\int 2 dt \)

OpenStudy (irishboy123):

\(v(2) = 16\)

OpenStudy (beggi9):

@Irishboy123 ah okay I see, thank you so much!